OpenStudy (anonymous):

Calculus Help!! Is there a faster way to do this? ∫csc^10tcot^3tdt

OpenStudy (anonymous):

Faster than what?

OpenStudy (anonymous):

XD that doesnt look good

OpenStudy (anonymous):

is it we can put u=csct and du=-csctcotdt? which is faster?

OpenStudy (anonymous):

No, I do not believe you can do u substitution in that case because of the powers on them.

OpenStudy (turingtest):

\[\int\csc^{10}t\cot^3tdt=\int\csc^{9}t\cot^2t(\csc t\cot t)dt\]\[=\int\csc^9t(1-\csc^2t)(\csc t\cot t)dt=\int\csc^9t-\csc^{11}t(\csc t\cot t)dt\]\[u=\csc t\]\[du=-\csc t\cot tdt\]

OpenStudy (turingtest):

whoops, second part is backwards...

OpenStudy (turingtest):

should have been\[\int\csc^9t(\csc^2t-1)(\csc t\cot t)dt=\int\csc^{11}t-\csc^9t(\csc t\cot t)dt\]now do the u sub above

OpenStudy (turingtest):

forgot parentheses above...\[\int\csc^9t(\csc^2t-1)(\csc t\cot t)dt=\int(\csc^{11}t-\csc^9t)(\csc t\cot t)dt\]

OpenStudy (anonymous):

thank you so much!!

OpenStudy (anonymous):

when do we have to use the reduction formulas?

OpenStudy (turingtest):

when all else fails I hate them personally, and almost always find a way around them sometimes I wind up integrating these types of things by parts (which is the same as the reduction formula) and don't even notice

OpenStudy (anonymous):

Thank you!!